OpenStudy (anonymous):

8. Caitlin had $402 in her bank account. She withdrew $15 each week to pay for a swimming lesson. She now has $237. a.    Write an equation that can be used to find the number of swimming lessons that she paid for.         b.    How many swimming lessons did she pay for?         c.    At the time she had $237, the cost of a lesson rose to $19. How many lessons can she pay for with her remaining $237?    

OpenStudy (tkhunny):

402 - 237 = ???

OpenStudy (anonymous):

a. 237 + 15p = 402, where p = number of lessons paid for b. 11 lessons c. 12 lessons is this right

OpenStudy (anonymous):

165

OpenStudy (anonymous):

@nader1

OpenStudy (tkhunny):

That will work. $165 is the total paid for the lessons. What is the number of lessons?

OpenStudy (anonymous):

11

OpenStudy (tkhunny):

11*15 = 165 There it is! Good work.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

@tkhunny medal please:)